Owl Shift: 6/7/99

Jason LaDue (jladue@thera.colorado.edu)
Mon, 7 Jun 1999 06:09:44 -0600 (MDT)

Shift Summary for Evening shift, 05/29/97

On Shift:
Jason LaDue
Elizabeth Turner

Other People Present: Ed Blucher and Rick Kessler.

Hours of beam delivered at all intensities: 5.167
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 3.167

Accounting of lost time:

Fresh Starting, No Beam Stop in, and debugging DA problem

========================================================================
Runs ended this shift:
13411 - MU_4E11
13412 - MU_4E11
13413 - MU_4E11

Tapes brought to FCC this shift: 0

========================================================================
Accesses during shift: None

========================================================================
Description of Shift:

00:02 - Left bend Quench
00:09 - Beam returns.
00:16 - SEM reaches ~6e11 for a couple of spills
02:38 - Controls problems, lose beam
02:52 - Beam returns
03:02 - Pipeline crate 2 is full
03:35 - Discover that Beam Stops were NOT in for the last half hour!
04:00 - Cannot Fresh Start
05:15 - BEAM_832 destination files have bad syntax, now using
BEAM_832_ECAL.
05:40 - Water leak in NS7. No water getting to NM2EU, and power
supply for NM3S is wet, estimate at least 2 hours.
06:13 - Hellium Pressure down to 80 PSI.

run spill time comment
------------------------------------------------------------------------
13411 00:58 Run Ended
13412 62 Beam Stops no longer in!
13412 03:28 Run Stopped
13413 05:22 Run Started,
13413 05:48 Correct triggers loaded for this run.